12) Solve the following two cases (the cases are independent). a.If you invest $5,000 today at 10% interest, what is the value of the investment at the end of 5 years? b.If you invest $1,200 at the end of each of the next 5 years and the investment earns 10% interest, what is the value of the investment at the end of 5 years?
Answer #1

a. Investment at the end of 5 year = $5000*1.6105 = $8052.55

b. Investment at the end of 5th year = 1200*6.1051 = $7326.12